OpenStudy (anonymous):

Find four consecutive odd integers where the product of the two smaller numbers is 64 less than the product of the two larger numbers

OpenStudy (shubhamsrg):

let nos. be (a-1),(a+1),(a+3),(a+5) given that : (a-1)(a+1) + 64= (a+3)(a+5) just simplify and solve for a..

OpenStudy (foolaroundmath):

taking (a-3), (a-1), (a+1), (a+3) as the four numbers would simplify calculations given: (a-3)(a-1) = (a+1)(a+3) - 64

OpenStudy (shubhamsrg):

i'd say same amount of effort would be req..